OpenStudy (anonymous):

f(x)=x^3-2x^2+x find all zeros by factoring each function

OpenStudy (anonymous):

need help

jhonyy9 (jhonyy9):

here you will get x(x^2 -2x +1) = x(x-1)^2

jhonyy9 (jhonyy9):

all zeros you can calculate now --- yes ?

OpenStudy (anonymous):

ok

jhonyy9 (jhonyy9):

push please the good answer

jhonyy9 (jhonyy9):

thank you

OpenStudy (anonymous):

you don't tell me your answer yet

jhonyy9 (jhonyy9):

all zeros are 0 and 1

jhonyy9 (jhonyy9):

so yes when x=0 and x=1

OpenStudy (anonymous):

why their answer is {0 , 1 mult.2}

OpenStudy (anonymous):

what mean of mult.2?

jhonyy9 (jhonyy9):

what is mult 2 ???

OpenStudy (anonymous):

I don't know their answer like that

jhonyy9 (jhonyy9):

the zeros are 0 and 1 because if you check your equation for this values of x like result you get zero

OpenStudy (anonymous):

but why ther write{o, 1 mult. 2} in their result

jhonyy9 (jhonyy9):

x(x-1)^2 so when x=0 this all equal zero when x=1 you get x(1-1)^2 what is zero too

jhonyy9 (jhonyy9):

who are their ?

OpenStudy (anonymous):

algebra 3review

OpenStudy (anonymous):

algebra 2 review

OpenStudy (anonymous):

their worksheet and answer

jhonyy9 (jhonyy9):

indifferent what their have wrote for this equation all zeros are for x=0 and for x=1

OpenStudy (anonymous):

maybe I think so

OpenStudy (anonymous):

ok thank for ur help.